### v4

# Retinex-RAWMamba: Bridging Demosaicing and Denoising for Low-Light RAW Image Enhancement

Abstract
Low-light image enhancement, particularly in cross-domain tasks such as mapping from the raw domain to the sRGB domain, remains a significant challenge. Many deep learning-based methods have been developed to address this issue and have shown promising results in recent years. However, single-stage methods, which attempt to unify the complex mapping across both domains, leading to limited denoising performance. In contrast, two-stage approaches typically decompose a raw image with color filter arrays (CFA) into a four-channel RGGB format before feeding it into a neural network. However, this strategy overlooks the critical role of demosaicing within the Image Signal Processing (ISP) pipeline, leading to color distortions under varying lighting conditions, especially in low-light scenarios. To address these issues, we design a novel Mamba scanning mechanism, called RAWMamba, to effectively handle raw images with different CFAs. Furthermore, we present a Retinex Decomposition Module (RDM) grounded in Retinex prior, which decouples illumination from reflectance to facilitate more effective denoising and automatic non-linear exposure correction. By bridging demosaicing and denoising, better raw image enhancement is achieved. Experimental evaluations conducted on public datasets SID and MCR demonstrate that our proposed RAWMamba achieves state-of-the-art performance on cross-domain mapping. The code is available at https://github.com/Cynicarlos/RetinexRawMamba.
1 Introduction
Existing deep learning methods, particularly those focused on low-light enhancement tasks, primarily operate in the sRGB domain. However, RAW images typically possess a higher bit depth than their RGB counterparts, meaning they retain a greater amount of original detail. Consequently, processing from RAW to RGB is often more effective. However, RAW and RGB are distinct domains with image processing algorithms tailored to their specific characteristics. For instance, in the RAW domain, algorithms prioritize denoising, whereas in the RGB domain, they focus on color correction. This difference often renders single-stage end-to-end methods [26, 12, 19] ineffective.
Figure 1: (a) A kind of demosaicing interpolation for RGGB Bayer Pattern and (b) the scanning in RAWMamba (black dashed line) and naive Mamba (purple dashed line). Note that only four directions of RAWMamba are drawn, reversing them gives four more directions, eight in all.
Demosaicing algorithms play a crucial role in converting RAW image to sRGB, with most traditional methods relying on proximity interpolation. Although some researchers have explored CNN-based approaches [42, 7] to map noisy RAW images to clean sRGB outputs, the limited receptive field inherent in convolutional networks often hampers their effectiveness in demosaicing tasks. To address this, Vision Transformers (ViTs) have been employed to expand the receptive field, but the attention mechanisms in ViTs are computationally intensive. The introduction of Mamba provides a more efficient balance between these trade-offs. However, existing Mamba scanning mechanisms do not adequately address the diverse characteristics of RAW images with different Color Filter Arrays (CFAs), highlighting the need for Mamba scanning methods specifically tailored to various CFAs.
Hence, we design a novel Mamba scanning mechanism for RAW format image(RAWMamba), which has a global receptive field and an attention mechanism with linear complexity that can better adapt to the data in this task. More importantly, as shown in Fig. 1 (b), naive Mamba scanning mechanism do not consider imaging properties, leading to limitations in feature extraction with CFA. In contrast, our RAWMamba introduces eight distinct scanning directions, fully accounting for all pixels in the immediate neighborhood of a given pixel while preserving the spatial continuity of the image. Specifically, the scanning directions encompass horizontal, vertical, oblique scanning from top left to bottom right, and oblique scanning from top right to bottom left. These four primary directions are mirrored to produce an additional four directions, resulting in a total of eight scanning directions.
Additionally, previous methods [5, 17] for processing short-exposure RAW images often rely on a simple linear multiplication of a prior for exposure correction. Specifically, short-exposure RAW images, which contain significant noise, are multiplied by the exposure time ratio of the corresponding long-exposure image. This approach assumes uniform exposure across the image, which is often unrealistic and can result in sub-optimal denoising and inaccurate brightness. By leveraging the success of the Retinex theory in low-light enhancement tasks for RGB images [21, 25, 33], we introduce a Retinex-based dual-domain auxiliary exposure correction method, namely Retinex Decomposition Module (RDM), which decouples illumination and reflection and realize automatic nonlinear exposure correction to achieve more efficient denoising effect and more accurate brightness correction. Furthermore, given the significant differences in noise distribution between different RAW domain and sRGB domain, we build upon the idea of decoupling the task into two sub-tasks: denoising on the raw domain [43, 18, 35] and cross-domain mapping [30, 46, 34, 16].
In general, we propose a Retinex-based decoupling network (Retinex-RAWMamba) for RAW domain denoising and low-light enhancement shown in Fig. 2. Our method decouples the tasks of denoising and demosaicing into two distinct sub-tasks, effectively mapping noisy RAW images to clean sRGB images. Specifically, for demosaicing sub-task, we introduce RAWMamba to fully consider all pixels in the immediate neighborhood of a certain pixel by utilizing eight direction mechanism. For the denoising sub-task, we propose the Retinex Decomposition Module, which enhances both denoising performance and brightness correction. Additionally, we introduce a dual-domain encoding stage enhance branch designed to leverage the meticulously preserved detail features from the raw domain, thereby compensating for the information loss that occurs during the denoising phase.
Our main contributions are summarized as follows:
  • We propose a Retinex-based decoupling Mamba network for RAW domain denoising and low-light enhancement (Retinex-RAWMamba). To our best knowledge, this is first attempt to introduce Mamba mechanism into low-light RAW image task.
  • We design a novel eight-direction Mamba scanning mechanism, to thoroughly account for the intrinsic properties of RAW images, and develop a Retinex Decomposition Module to bridging denoising capabilities and exposure correction.
  • We evaluate the proposed method on two benchmark datasets quantitatively and qualitatively. The comprehensive experiments show that the proposed method outperforms other state-of-the-art methods in PSNR, SSIM and LPIPS with a comparable number of parameters.
Figure 2: The overall architecture of our proposed Retinex-RAWMamba and (a) Retinex Decomposition Module, (b) Simple Denoising Block and (c) Domain Adaptive Fusion
2 Related Work
2.1 Low Light Enhancement on Raw Domain
In the Raw domain low-light enhancement task, researchers have proposed some innovative approaches. Since the task can be split into two sub-tasks, RAW domain denoising and color correction, some of the work only focuses on one of the sub-tasks. For example, on the raw domain denoising task, there are noise modeling with deep learning methods [2, 4, 8, 9, 39, 51], which ultimately compute evaluation metrics on the raw domain. After the release of the SID public dataset by Chen at al.[5] 2018, researchers have proposed many works that address both tasks simultaneously. These works can be further categorized into single-stage approaches and multi-stage approaches. Single-stage methods [5, 45] aims to map noisy raw to clean sRGB by training a single model. For instance, SID [5] only used a simple UNet to accomplish this task. DID [26] proposed a deep neural network based on residual learning for end-to-end extreme low-light image denoising. SGN [12] introduced a self-guided network, which adopted a top-down self-guidance architecture to better exploit image multi-scale information. Since the ISP undergoes many nonlinear transformations, it is still difficult to learn for a single neural network, and it can only be realized by piling up a large number of parameters, which leads to inefficiencies, and thus multi-stage methods came into being. Multi-stage methods [15, 42, 50] achieve better results by decoupling the tasks, this idea effectively reduces the ambiguity between different domains. For instance, Huang et al. [15] proposed intermediate supervision on the raw domain, while Dong et al. [7] did that on the monochrome domain. DNF [17] introduced a decoupled two-stage net with weight weight-shared encoder to reduce the number of parameters while achieving good results.
2.2 Mamba in Vision Task
State Space Models (SSM) are recently introduced to deep learning since they can effectively model long range dependencies. For instance, [11] proposes a Structured State-Space Sequence (S4) model and recently, [10] proposes Mamba, which outperforms Transformers at various sizes on large-scale real data and enjoys linear scaling in sequence length. In addition to Mamba’s great work on NLP tasks, researchers have also made many attempts and achieved good results on visual tasks, such as classification [6, 44], segmentation [24, 37, 41, 29, 22], generation [31, 14], and image restoration [13, 32, 48, 49, 27, 1]. EfficientVMamba [27] presents the Efficient 2D Scanning (ES2D) method, utilizing atrous sampling of patches on the feature map to speed up training. VMamba [23] incorporates a Cross-Scan Module (CSM), which converts the input image into sequences of patches along the horizontal and vertical axes, and it enables the scanning of sequences in four distinct directions. That is, each pixel integrates information from the four surrounding pixels. FreqMamba [48] introduces complementary triple interaction structures including spatial Mamba, frequency band Mamba, and Fourier global modeling, which utilizes the complementarity between Mamba and frequency analysis for image deraining.
3 Method
3.1 Preliminaries
3.1.1 State Space Model (SSM)
SSM is a linear time-invariant system that maps input $x(t)\in{\mathbb{R}}^{L}$ to output $y(t)\in{\mathbb{R}}^{L}$. SSM can be formally represented by a linear ordinary differential equation (ODE),
$h^{\prime}(t)$ $=\mathbf{A}h(t)+\mathbf{B}x(t),$ (1)
$y(t)$ $=\mathbf{C}h(t)+\mathbf{D}x(t)$
SSM is continuous-time model, presenting significant challenges when integrated into deep learning algorithms. To address this issue, discretization becomes a crucial step. Denote $\Delta$ as the timescale parameter. The zero-order hold (ZOH) rule is usually used for discretization to convert continuous parameters $\mathbf{A}$ and $\mathbf{B}$ in Eq. 1 into discrete parameters $\overline{\mathbf{A}}$ and $\overline{\mathbf{B}}$. Its definition is as follows:
$\overline{\mathbf{A}}$ $=exp(\Delta\mathbf{A}),$ (2)
$\overline{\mathbf{B}}$ $=(\Delta\mathbf{A})^{-1}(exp(\Delta\mathbf{A})-\mathbf{I})\cdot\Delta\mathbf{B}$
After the discretization of A, B, the discretized version of Eq. 1 using a step size $\Delta$ can be rewritten as:
$h_{k}$ $=\overline{\mathbf{A}}h_{k-1}+\overline{\mathbf{B}}x_{k},$ (3)
$y_{k}$ $=\mathbf{C}h_{k}+\mathbf{D}x_{k}$
Finally, the models compute output through a global convolution as following:
$\overline{\mathbf{K}}$ $=(\mathbf{C}\overline{\textbf{B}},\mathbf{C}\overline{\mathbf{AB}},...,\mathbf{C}{\overline{\mathbf{A}}}^{L-1}\overline{\mathbf{B}})$ (4)
$\mathbf{y}$ $=\mathbf{x}\ast\overline{\mathbf{K}}$
where L is the length of the input sequence x, and $\overline{\mathbf{K}}\in{\mathbb{R}}^{L}$ is a structured convolutional kernel.
3.2 Overall Pipeline
The overall pipeline is shown in Fig. 2. First, we preprocess the low-exposure noisy single-channel raw image by multiplying it with the exposure time ratio of the long-exposure ground truth (GT). Then, based on the Color Filter Array (CFA) pattern, we pack it into a multi-channels input. Specifically, for Bayer format, we pack the input $\textbf{X}\in\mathbb{R}^{H\times W\times 1}$ into four channels input $\textbf{X}_{packed}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 4}$; for XTrans format, we pack the input into nine channels input $\textbf{X}_{packed}\in\mathbb{R}^{\frac{H}{3}\times\frac{W}{3}\times 9}$. Both stages of Retinex-RAWMamba are built upon the UNet-based [28] encoder-decoder architecture. The first stage of the overall framework is dedicated to raw domain denoising. Initially, The Retinex Decomposition Module (RDM) processes the input to generate two feature maps L and R, L will be multiplied by the original input to obtain $\textbf{X}_{in}$ and R will be used later. $\textbf{X}_{in}$ will pass the denoising stage and generate the first output $\textbf{O}_{1}\in\mathbb{R}^{H\times W\times C_{in}}$. Then $\textbf{X}_{in}$ will pass the demosaicing stage to generate the second output $\textbf{O}_{2}\in\mathbb{R}^{H\times W\times 3}$. The overall loss function is then calculated against both ground truth RAW and ground truth RGB images, providing the supervision signal for both domains and guiding the optimization of whole model.
Figure 3: Details of (a) RAWMamba, (b) RAWSSM and (c) SS2D
3.3 RAWMamba
The details of RAWMamba and is shown in Fig. 3 (a). The RAWSSM leverages the naive visual mamba in MambaIR [13], with an innovative scanning mechanism. In the ISP process from Raw to RGB, proximity interpolation is commonly employed for demosaicing and often involves considering all eight closely connected locations around a given position, and Fig. 1 (a) gives an example with Bayer pattern raw image, (b) shows the scanning in RAWMamba (black dashed line) and naive Mamba (purple dashed line). The naive scanning method fails to consider the continuity of scanning, resulting in a lack of continuity between the end of each row/column and its bottom/right side. This leads to gaps in image semantics, which hinders image reconstruction. To address this issue, we propose using a Z-scan. That is when the scan reaches the end of each row/column, the reverse scan starts from the next row/column immediately adjacent to the last pixel. However, there are still limitations with this scanning method as it does not take into account all eight surrounding pixels when certain pixels are close to each other at the top, bottom, left, and right positions. Taking into consideration the characteristics of this task, we introduce Eight direction Mamba.
The detail of the our proposed scan mechanism is shown in Fig. 3 (c). Specifically, for a feature map $\textbf{{F}}\in{\mathbb{R}}^{C\times H\times W}$, we first flip its even rows (fer) and columns (fec) to get $\textbf{{F}}_{fer}\in{\mathbb{R}}^{C\times H\times W}$ and $\textbf{{F}}_{fec}\in{\mathbb{R}}^{C\times H\times W}$, respectively. Then we flatten $\textbf{{F}}_{fer}$ and $\textbf{{F}}_{fec}$ to get the first two directions’ scanning $\textbf{{F}}_{1}\in{\mathbb{R}}^{C\times HW}$ and $\textbf{{F}}_{2}\in{\mathbb{R}}^{C\times HW}$. Then we can get the feature of the oblique scan $\textbf{{F}}_{3}$ as following:
$(r,c)=(ceil(\frac{H\times W}{W+1}),(W+1))$ (5)
$\textbf{{F}}_{pad}=pad(flatten(\textbf{{F}}),(C,r\times c))$
$\textbf{{F}}_{reshaped}=reshape(\textbf{{F}}_{pad},(C,r,c))$
$\textbf{{F}}_{select}=ms(trans(\textbf{{F}}_{reshaped}),trans(\textbf{{mask}}))$
$\textbf{{F}}_{3}=reshape(\textbf{{F}}_{select},(C,H\times W))$
where $pad(\textbf{{X}},shape)$ is a function to pad X to new shape, mask and $\textbf{{F}}_{reshaped}$ have the same shape, and the first $H\times W$ elements of mask are true and the rest are false, trans(X) is a function to transpose $\textbf{{X}}\in{\mathbb{R}}^{C\times H\times W}$ to $\textbf{{X}}_{trans}\in{\mathbb{R}}^{C\times W\times H}$, $ms(\textbf{{X}},\textbf{{mask}})$ is a function to select elements from X based on the position where mask is true. The other oblique scan feature $\textbf{{F}}_{4}$ can be obtained similarly, and then we can invert these four features to get another four directions features, eight in total, which is $\{\textbf{{F}}_{i}\in{\mathbb{R}}^{C\times HW},i=1,2,...8\}$. At this point, the scanning of the eight directions is completed. And after the SSM, we get $\{\overline{\textbf{{F}}}_{i}\in{\mathbb{R}}^{C\times HW},i=1,2,...8\}$, we then merge them by adding them up and reshape these eight features to the original shape to get a single feature, that is,
$SS2D(\textbf{{F}})=Reshape(\sum_{i=1}^{8}\overline{\textbf{{F}}}_{i},(C,H,W))$ (6)
For the RAWSSM, given an input X, it can be formulated as follows:
$x,z$ $=chunk(Linear(\textbf{{X}}))$ (7)
$x$ $=LN(SS2D(SiLU(Conv_{3}(x))))$
$out$ $=x*SiLU(z)$
where out is the output of RAWSSM, LN is layer normalization, $Conv_{3}$ is the convolution operation with a kernel size of $3\times 3$, SiLU is the activate function.
And for the proposed RAWMamba, given an input X, it can be simply formulated as:
$t$ $=\alpha\textbf{{X}}+RAWSSM(LN(\textbf{{X}})))$ (8)
$out$ $=\beta t+CA(GELU(Conv(LN(t))))$
where, out is the output of RAWMamba, $\alpha$ and $\beta$ are parameters that can be learned, CA is channel attention.
3.4 Retinex Decomposition
Low-light enhancement methods based on retinex theory have been successful in RGB domain [38, 40, 3], so we propose dual-domain Retinex Decomposition Module. RDM can decompose image $\textbf{{X}}\in{\mathbb{R}}^{H\times W\times C_{in}}$ into the reflection component $\textbf{{R}}\in{\mathbb{R}}^{H\times W\times C}$ and the illumination component $\textbf{{L}}\in{\mathbb{R}}^{H\times W\times C_{in}}$. The details of RDM are shown in Fig. 2 (a). The module first takes the average value of the input image $\textbf{{X}}\in{\mathbb{R}}^{H\times W\times C_{in}}$ in the channel dimension to obtain $\textbf{{M}}\in{\mathbb{R}}^{H\times W\times 1}$, concatenates it in the channel dimension, and then passes several convolutions and a GELU activate function to obtain the first output $\textbf{{R}}\in{\mathbb{R}}^{H\times W\times C}$, and then passes a $1\times 1$ convolution to obtain the light map L , which will be multiplied by the original input to pre-adjust the light. Specifically,
L $=Convs_{1,5,3}\{cat[\textbf{{X}},mean(\textbf{{X}},dim=-1)]\}$ (9)
R $=Conv_{1}(\textbf{{L}})$
where cat refers to the concatenation of two feature maps on the channel dimension, $Convs_{1,5,3}$ donates a series of convolution with kernel size 1, 5 and 3.
Figure 4: The visualization results between our method and the state-of-the-art methods (Zoom-in for best view).
3.5 Dual-domain Encoding Stage Enhance Branch
Considering that the feature R obtained from the RDM contains most of the details that could be lost after the first stage, we make full use of these features in both domains and in order to reduce the amount of calculation, we propose dual-domain encoding stage enhance branch, which does not be used in the decoding stages. Specifically, after obtaining R, we will simply downsample it to get four feature maps at each layer, which are donated as $\textbf{{R}}_{i},i=1,2,3,4\}$, $\textbf{{R}}_{i}$ is the $i_{th}$ layer light feature that will be fused later with DAF for auxiliary automatic exposure correction at layer $i$. At the $i_{th}$ layer of the encoding stage during the denoising phase, the denoising feature $\textbf{{dn}}_{i}$ will firstly fused with $\textbf{{R}}_{i}$ before passing the SDB. Similarly, at the $i_{th}$ layer of the encoding stage during the demosaicing phase. the the demosaicing feature $\textbf{{dm}}_{i}$ will firstly fused with $\textbf{{R}}_{i}$ before passing the RAWMamba. This method of performing fusion only in the encoding stage fully utilizes the features that are not lost to improve the ability to restore details while reducing the amount of calculation.
3.6 Domain Adaptive Fusion
The details of DAF are shown in Fig. 2 (c), previous feature map will be firstly concatenated with current feature map at the same level, and this result will be multiplied with previous feature map after the convolution, then it will pass through a convolution with a residual addition. And we can get the fused feature map after a final convolution. Specifically, for the two feature maps $\textbf{{pre}}\in{\mathbb{R}}^{H\times W\times C}$ and $\textbf{{cur}}\in{\mathbb{R}}^{H\times W\times C}$, they will be fused as following:
$\textbf{{T}}=Conv_{3}(cat(\textbf{{pre}},\textbf{{cur}}))$ (10)
$=Conv_{1}(CA(\textbf{{T}}))$ T ) )
$\textbf{{T}}=\textbf{T}\odot Conv_{1}(GELU(\textbf{{pre}}))$
$=Conv_{1}(GELU(\textbf{{T}}))$ T ) )
$Out(\textbf{{cur}},\textbf{{pre}})=Conv_{1}(\textbf{{T}}+\textbf{{cur}})$
3.7 Loss Function
Traditional low-level vision tasks generally use L1 Loss, and we also follow that, but our task involves different sub-tasks on two domains, Raw domain and sRGB domain, so the loss can be expressed as following:
$L_{total}$ $=\alpha L_{raw}+\beta L_{srgb}$
$=\alpha{||\hat{Y}_{raw}-GT_{raw}||}_{1}+\beta{||\hat{Y}_{srgb}-GT_{srgb}||}_{1}$
where $Y_{raw}$ is the raw image after denoised, $Y_{srgb}$ is the sRGB image after the second stage, $GT_{srgb}$ is the sRGB image obtained from raw ground truth after post-processing by Rawpy as previous work did. And $\alpha$ and $\beta$ defaults to 1.0 in our experiments.
Table 1: Quantitative results of RAW-based LLIE methods on the Sony and Fuji subsets of SID. The top-performing result is highlighted in red, while the second-best is shown in green. Metrics marked with $\uparrow$ indicate that a higher value is better, and those marked with $\downarrow$ indicate that a lower value is better. ‘-’ indicates the result is not available.
Category Method Venue #Params.(M) Sony Fuji
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Single-Stage SID CVPR2018 7.7 28.96 0.787 0.356 26.66 0.709 0.432
DID ICME2019 2.5 29.16 0.785 0.368 - - -
SGN ICCV2019 19.2 29.28 0.790 0.370 27.41 0.720 0.430
LLPackNet BMVC2020 1.2 27.83 0.755 0.541 - - -
RRT CVPR2021 0.8 28.66 0.790 0.397 26.94 0.712 0.446
Multi-Stage EEMEFN AAAI2020 40.7 29.60 0.795 0.350 27.38 0.723 0.414
LDC CVPR2020 8.6 29.56 0.799 0.359 27.18 0.703 0.446
MCR CVPR2022 15.0 29.65 0.797 0.348 - - -
RRENet TIP2022 15.5 29.17 0.792 0.360 27.29 0.720 0.421
DNF CVPR2023 2.8 30.62 0.797 0.343 28.71 0.726 0.391
Ours - 6.2 30.76 0.810 0.328 29.02 0.743 0.382
4 Experiments
4.1 Datasets and Experiments Environments
4.1.1 SID Dataset
For Sony subset, there are totally 1865 raw image pairs in the training set. Each pair of images contains a short exposure and a long exposure, the short exposure is used as noisy raw, and the long exposure is used as $GT_{raw}$. The original size of all images is $2848\times 4256$. Limited by GPU memory, the data is preprocessed before training, first pack into $4\times 1424\times 2128$, then randomly crop a patch with shape $4\times 512\times 512$ as the input with random data augmentation, such as horizontal/vertical flipping. For the test set, we referred to the DNF[17] settings and deleted the three misaligned scene images.
For Fuji subset, similar to Sony subset, 1655 and 524 raw image pairs for training and testing, respectively. The original size of it is $4032\times 6032$, since its CFA (Color Filter Array) is X-Trans instead of Bayer, we pack it into $9\times 1344\times 2010$ and randomly crop a patch with shape $9\times 384\times 384$ as the input.
4.1.2 MCR Dataset
The MCR [7] dataset contains 4980 images with a resolution of $1280\times 1024$, including 3984 low-light RAW images, 498 monochrome images (not be used for us) and 498 sRGB images. With indoor and outdoor scenes, different exposure times are set, 1/256s to 3/8s for indoor scenes and 1/4096s to 1/32s for outdoor scenes. And we obtained the raw ground truth as DNF [17] did. The preprocessing is similar to SID dataset, but we don’t randomly crop a patch as the input.
4.1.3 Implementation Details
During training, the batch size is 1 and the initial learning rate is 1e-4, and we use the cosine annealing strategy to reduce it to 1e-5 at the 200th epoch. Adamw optimizer is used and the betas parameter is [0.9,0.999] and the momentum is 0.9. The training and testing is completed by a NVIDIA 3090 (24G), A40 (48G),respectively due to the limitation of GPU memory. We also provide the code of merging test on a 24G GPU. Note that the results of merging test are a litter bit smaller than that testing with whole image. And we use PSNR, SSIM [36] and LPIPS [47] as the quantitative evaluation metrics.
4.2 Comparison with State-of-the-Arts
We conduct experiments on SID [5] dataset including Sony and Fuji subsets and MCR [7] dataset, and compare with previous SOTA methods including SID [5], DID [26], SGN [12], EEMEFN [50], LDC [42], LLPackNet [20], RRT [19], MCR [7], RRENet [15] and DNF [17].
The results are presented in Tab. 1 and 2. As observed, most single-stage methods underperform compared to multi-stage methods, demonstrating the feasibility and effectiveness of the multi-stage approach for noisy RAW to clean sRGB cross-domain mapping. On the SID dataset, our proposed method outperforms all metrics among multi-stage approaches, while maintaining a smaller parameter count. Specifically, on the Sony and Fuji subsets, our method achieves a PSNR increase of 0.14 dB and 0.31 dB, respectively, an SSIM improvement of 0.011 and 0.017, and an LPIPS reduction of 0.015 and 0.009, compared to the best existing method.
For the MCR dataset, as shown in Tab. 2, while our improvement in SSIM is modest, we achieve a significant PSNR increase of 1.14 dB, a 3.6% enhancement over the second-best method. Additionally, we selected several previous state-of-the-art (SOTA) methods and visualized their performance on the SID Sony dataset, as shown in Fig. 4. Three scenarios are depicted, each containing two sub-regions. In the first two scenes, most other methods produce a green tint to the image. In the third scene, these methods often fail to preserve details adequately. In contrast, our proposed method closely aligns with the ground truth in both color and detail, effectively achieving denoising and color enhancement in the raw domain under low-light conditions.
4.3 Ablation Studies
Table 2: Quantitative results on MCR [7] dataset. The top-performing result is highlighted in red, while the second-best is shown in green. Metrics marked with $\uparrow$ indicate that a higher value is better, and those marked with $\downarrow$ indicate that a lower value is better.
Category Method PSNR $\uparrow$ SSIM $\uparrow$
Single-Stage SID 29.00 0.906
DID 26.16 0.888
SGN 26.29 0.882
RRT 25.74 0.851
Multi-Stage LDC 29.36 0.904
MCR 31.69 0.908
DNF 32.00 0.915
Ours 33.14 0.914
To demonstrate the validity of our proposed method, we perform ablation experiments on the SID Sony dataset. We first propose a baseline model that consists only of SDB and the unmodified naive visual mamba in MambaIR [13] and GFM in DNF [17]. Tab. 3 shows the results of adding or replacing the corresponding module based on the baseline, where RAWM stands for replacing naive Mamba with RAWMamba, RDM stands for adding RDM module, and DAF stands for replacing GFM with DAF module. All the ablation experiments were conducted in the same environment.
First, we replaced the baseline’s naive Mamba with the proposed RAWMamba. The results showed increases of 0.41 dB in PSNR and 0.012 in SSIM, demonstrating that our RAWMamba, with its eight-directional scanning mechanism, performs well in the demosaicing task. Next, we incorporated the proposed RDM for denoising and automatic exposure correction. The results indicated that although SSIM did not improve, PSNR increased by an additional 0.27 dB. This suggests that the initial exposure of the images was indeed problematic, and our RDM effectively enhances denoising and exposure correction. Finally, we replaced all GFM components in the network with our proposed DAF to improve the stability of the training process. This led to further gains, with PSNR and SSIM increasing by 0.06 dB and 0.001, respectively.
Furthermore, we conducted a straightforward visualization of the ablation study, depicted in Fig. 5. The incorporation of RAWMamba into the baseline model effectively reduces noise and mitigates the green color distortion. Comparatively, the Retinex-RawMamba approach demonstrates superior color correction capabilities and attains the highest PSNR and SSIM scores. This clearly indicates that our proposed method outperforms others in terms of both detail preservation and color accuracy.
Table 3: Ablation study on SID Sony dataset.
Baseline RAWM RDM DAF PSNR SSIM
30.04 0.797
30.45 0.809
30.70 0.809
30.76 0.810
Figure 5: The visualization results for ablation studies (Zoom-in for best view).
5 Conclusion
For the task of denoising and enhancing RAW images under low-light conditions, we introduce Retinex-RAWMamba, a novel two-stage cross-domain network. Our approach extends the capabilities of the traditional Vision Mamba by incorporating RAWMamba, which exploits the inherent properties of demosaicing algorithms in ISP to achieve enhanced color correction and detail retention. Additionally, we integrate Retinex theory through our Retinex Decomposition Module, facilitating automatic exposure correction and yielding RGB images with improved illumination and brightness fidelity. Comprehensive theoretical analysis and experimental validation underscore the effectiveness and significant potential of our method.